# Convergence of Fast Policy Iteration in Markov Games and Robust MDPs

### v1

@@ 8 $\bm{v}^{k}\leftarrow\bm{v}^{k-1}+\beta^{i_{k}}\cdot\bm{d}^{k}$;
To derive the FT algorithm, one interprets PAI as the pure Newton’s method for solving $\min_{\bm{v}\in\mathbb{R}^{S}}\psi_{2}(\bm{v})^{2}$ (Filar and Tolwinski 1991; Filar and Vrieze 1997). Recall that the pure Newton’s method direction $\bm{d}^{k}$ in iteration $k\in\mathbb{N}$ is
$\bm{d}^{k}$ $:=-(\nabla^{2}\psi_{2}(\bm{v}^{k})^{2})^{-1}\nabla\psi_{2}(\bm{v}^{k})^{2}$
$=\left(\bm{I}-\gamma\bm{P}^{\bm{\pi}^{k+1},\bm{\sigma}^{k+1}}\right)^{-1}\left(\mathfrak{T}^{\star}\bm{v}^{k}-\bm{v}^{k}\right).$
@@
Theorem 3.3 in (Filar and Tolwinski 1991) states that Algorithm˜1 is guaranteed to converge to the optimal value function. However, there is a gap in the proof. In particular, while each step of the iteration reduces the value function, it is not guaranteed that a step size satisfying Armijo’s rule exists. Since the gradient of $\bm{v}\mapsto\psi_{2}(\bm{v})^{2}$ may be discontinuous, it is possible that no $i$ in Line 1 in Algorithm˜1 satisfies the inequality; leading to an infinite loop in the search for the step size. We construct a simple MDP example demonstrating this behavior to show that this can happen.
Figure 1: Plot of $\psi_{2}(\bm{v})^{2}$ projected onto the plane that spans the initial value function, optimal value function, and the step direction share.
@@
We note that Filar and Tolwinski (1991, theorem 3.3) assumes that FT is initialized to a constant value determined by the maximum reward instead of a zero vector. However, the algorithm makes no progress even with such initialization as Remark˜B.1 in the appendix demonstrates.
We now discuss the gap in the proof of convergence in Filar and Tolwinski (1991, theorem 3.3) which Theorem˜4.2 contradicts.As noted in Filar and Tolwinski (1991, theorem 2.1) the function $\bm{v}\mapsto\psi_{2}(\bm{v})^{2}$ is differentiable almost everywhere. However, because the function is not differentiable everywhere, Armijo’s rule fails to find a positive step size. Example˜4.1 initializes FT in exactly a point of non-differentiability. One attempt to circumvent this problem would be to argue that the probability of being at a point of non-differentiability is zero. However, it may be possible to modify our example so that the initialization does not happen at a point of non-differentiability. Yet, the line search method will take increasingly smaller steps, such that it approaches the point of non-differentiability without ever passing it. Because it is unclear how one may rectify the non-differentiability to ensure Newton’s method’s convergence, we propose an alternative approach in the following section.
@@ 10    $\bm{v}^{k}\leftarrow\bm{u}^{k,l}$ ;
RCPI, summarized in Algorithm˜2, can be viewed as a direct modification of FT in Algorithm˜1. The first two steps of RCPI’s iteration are identical to FT. First, RCPI jointly updates both the primary and adversarial policies to be greedy with respect to the current value function. Second, RCPI evaluates the value function for the updated policies. Simply adopting this value function would lead to PAI (Algorithm˜3 in the appendix), which is prone to getting stuck in infinite cycles (Van der Wal 1978). Such infinite cycles must involve steps that do not decrease the residual. RCPI detects when the residual does not decrease sufficiently and reverts to a value function update to guarantee its reduction. As a result, RCPI will never cycle or terminate before reaching the optimal value function.
@@
We now turn to the proof of RCPI’s correctness and computation complexity. First, we need to discuss the worst-case runtime of the Bellman backups. For s-rectangular $L_{1}$ robust MDPs the runtime of computing $\mathfrak{T}^{\delta}\bm{v}$ and $\mathfrak{B}^{\delta}\bm{v}$ is (Ho, Petrik, and Wiesemann 2021)
@@ 6 Numerical Results
To evaluate the effectiveness of RCPI a series of examples were solved using a range of algorithms including PAI, Hoffman Karp (HK), Filar Tolwinski’s algorithm (FT), robust value iteration (VI), a variation of Hoffman Karp (PPI) (Ho, Petrik, and Wiesemann 2021), a variation on PAI (WS) (Winnicki and Srikant 2023), and our algorithm (RCPI). To simplify comparison, RCPI’s hyperparameter $m$ was either set to 0, producing a method that never fixed its evaluation step, called RCPI0, or $m$ was left unbounded, making a method that always fixed its evaluation step, called RCPI.
For each domain, there was a smaller set of problems solved with $\gamma$ set to $0.5$, $0.75$, $0.9$, and $0.99$, and a larger set of problems where we set $\gamma$ to $0.9$, $0.99$, and $0.999$. We use the smaller problems in order to evaluate the slower algorithms in a reasonable time. The larger problems can only be solved by the faster methods within our time limits. We allocated a larger time budget to VI to establish a comprehensive reference.
@@
Several domains were tested for robust MDPs including gamblers ruin (Kallenberg 2022, 1.3.1), gridworld (Sutton and Barto 2018, Example 6.5), and inventory management (Puterman 2005, 3.2.1). From Table˜1, each algorithm maintained the same relative performance from Markov games, except that WS and VI did comparatively better in the gambler’s ruin. WS and VI did well because the optimal betting scheme in the non-robust version of gambler’s ruin is to bet \$1 if the win rate is greater than $50\%$ and to bet all money otherwise. The gambler repeats this action until reaching the maximum capital, when it obtains the reward. The result is a singular optimal state trajectory following the current state. The reward from obtaining the maximum capital does not affect the current state’s policy until there is a Bellman backup for every state in the optimal trajectory following it. This type of domain favors methods with inexpensive evaluations, as only states that reward has been reached via policy improvement will be worth evaluating. Time spent doing exact evaluations for the other states does not provide any benefit.
@@ 7 Conclusion
It remains to be seen if there is a way to fix the algorithm of Filar and Tolwinski that keeps the Newtonian interpretation of PAI with $\psi_{2}(\bm{v})^{2}$ as the objective function. Such a solution could provide insight into how to deal with discontinuities caused by the min and max operations more generally. It is not known how to optimize the hyperparameter of RCPI $m$ which could reduce the level of knowledge required to use RCPI effectively. We suspect that there is no optimal selection of $m$ and that increasing it has the potential to reduce the expected solve time while increasing the runtime upperbound.
@@ A.2 Definitions for Robust MDPs
In this section, we summarize the pertinent definitions of value functions and Bellman operators in Robust MDPs. The value function $\bm{v}^{\bm{\pi},\bm{p}}\in\mathbb{R}^{S}$ is defined as for each $\bm{\pi}\in\Pi$ and $\bm{p}\in\mathcal{P}$ as (Wiesemann, Kuhn, and Rustem 2013; Ho, Petrik, and Wiesemann 2022):
$v^{\bm{\pi},\bm{p}}_{s}\;:=\;\mathbb{E}_{\bm{\pi},\bm{p}}^{s}\left[\sum_{t=0}^{\infty}\gamma^{t}r(\tilde{s}_{t},\tilde{a}_{t},\tilde{b}_{t},\tilde{s}_{t+1})\right],\quad\forall s\in\mathcal{S}.$ (14)
@@
$P^{\bm{\pi},\bm{p}}_{s,s^{\prime}}:=\sum_{a\in\mathcal{A}_{s}}\pi_{a}(s)\cdot p_{s^{\prime}}(s,a),\qquad r^{\bm{\pi}}_{s}:=\sum_{a\in\mathcal{A}_{s}}\pi_{a}(s)\cdot r(s,a).$
@@
The Bellman operators $\mathfrak{T}^{\bm{\pi},\bm{p}}$ and $\mathfrak{T}^{\star}$ are monotone and $\gamma$-contractive in the $L_{\infty}$ norm (Kallenberg 2022, theorem 10.5). They can be computed by solving a linear program (Ho, Petrik, and Wiesemann 2022, appendix C) or by a specialized algorithm (Ho, Petrik, and Wiesemann 2022, section 6).
@@ 4   Select $(\bm{\pi}^{k},\bm{\sigma}^{k})\in\mathfrak{B}^{\star}\bm{v}^{k-1}$;
Consider the policy such that $\bm{\sigma}_{1}(s_{1})=[1,0]$. Then, Algorithm˜1 computes the step direction as:
@@
To satisfy the starting constraint given from (Filar and Tolwinski 1991, theorem 3.3) that $\bm{v}^{0}=r_{\max}\cdot\bm{1}$ where $r_{\max}=\max_{a\in\mathcal{A},b\in\mathcal{B},s^{\prime}\in\mathcal{S}}|r(a,b,s^{\prime})|$, consider a MG with the transition probabilities and rewards defined in Figure˜4 and the discount factor $\gamma=0.8$.
@@
which is always positive for $\alpha\in(0,1]$, thus using the initial value from (Filar and Tolwinski 1991, theorem 3.3) does not fix FT’s convergence issues.
@@ Appendix C Additional Material for Section 5
This linear program contains $A+B$ constraints and $A$ variables, requiring $O\left((A+B)^{1.5}A^{2}\log(\delta^{-1})\right)$ arithmetic operations to solve to a precision of $\delta$ (Nemirovski 2004, section 10.1). Therefore, the total cost of computing $\mathfrak{B}^{\delta}\bm{v}$ and $\mathfrak{T}^{\delta}\bm{v}$ is the same as generating $\bm{G}^{\bm{v},s}$ and solving the linear program (C.1) for every state, producing the bound in (9). ∎

### v2

@@ 8 $\bm{v}^{k}\leftarrow\bm{v}^{k-1}+\beta^{i_{k}}\cdot\bm{d}^{k}$;
To derive the FT algorithm, one interprets PAI as the pure Newton’s method for solving $\min_{\bm{v}\in\mathbb{R}^{S}}\psi_{2}(\bm{v})^{2}$ (Filar1991; Filar1996). Recall that the pure Newton’s method direction $\bm{d}^{k}$ in iteration $k\in\mathbb{N}$ is
$\bm{d}^{k}$ $:=-(\nabla^{2}\psi_{2}(\bm{v}^{k-1})^{2})^{-1}\nabla\psi_{2}(\bm{v}^{k-1})^{2}$
$=\left(\bm{I}-\gamma\bm{P}^{\bm{\pi}^{k},\bm{\sigma}^{k}}\right)^{-1}\left(\mathfrak{T}^{\star}\bm{v}^{k-1}-\bm{v}^{k-1}\right).$
@@
Theorem 3.3 in (Filar1991) states that Algorithm˜1 is guaranteed to converge to the optimal value function. However, there is a gap in the proof. In particular, while each step of the iteration reduces the value function, it is not guaranteed that a step size satisfying Armijo’s rule exists. Since the gradient of $\bm{v}\mapsto\psi_{2}(\bm{v})^{2}$ may be discontinuous, it is possible that no $i$ in Line 1 in Algorithm˜1 satisfies the inequality; leading to an infinite loop in the search for the step size. We construct a simple MDP example demonstrating this behavior to show that this can happen.
Figure 1: Plot of $\psi_{2}(\bm{v})^{2}$ projected onto the plane that spans the initial value function, optimal value function, and the step direction.
@@
We note that Filar1991 assumes that FT is initialized to a constant value determined by the maximum reward instead of a zero vector. However, the algorithm makes no progress even with such initialization as shown in the appendix.
We now discuss the gap in the proof of convergence in Filar1991 which Theorem˜4.2 contradicts. As noted in Filar1991 the function $\bm{v}\mapsto\psi_{2}(\bm{v})^{2}$ is differentiable almost everywhere. However, because the function is not differentiable everywhere, Armijo’s rule fails to find a positive step size. Example˜4.1 initializes FT in exactly a point of non-differentiability. One attempt to circumvent this problem would be to argue that the probability of being at a point of non-differentiability is zero. However, it may be possible to modify our example so that the initialization does not happen at a point of non-differentiability. Yet, the line search method will take increasingly smaller steps, such that it approaches the point of non-differentiability without ever passing it. Because it is unclear how one may rectify the non-differentiability to ensure Newton’s method’s convergence, we propose an alternative approach in the following section.
@@ 10    $\bm{v}^{k}\leftarrow\bm{u}^{k,l}$ ;
RCPI, summarized in Algorithm˜2, can be viewed as a direct modification of FT in Algorithm˜1. The first two steps of RCPI’s iteration are identical to FT. First, RCPI jointly updates both the primary and adversarial policies to be greedy with respect to the current value function. Second, RCPI evaluates the value function for the updated policies. Simply adopting this value function would lead to PAI (see the appendix), which is prone to getting stuck in infinite cycles (VanderWal1978). Such infinite cycles must involve steps that do not decrease the residual. RCPI detects when the residual does not decrease sufficiently and reverts to a value function update to guarantee its reduction. As a result, RCPI will never cycle or terminate before reaching the optimal value function.
@@
We now turn to the proof of RCPI’s correctness and computation complexity. First, we need to discuss the worst-case runtime of the Bellman backups. For s-rectangular $L_{1}$ robust MDPs the runtime of computing $\mathfrak{T}^{\delta}\bm{v}$ and $\mathfrak{B}^{\delta}\bm{v}$ is (Ho2021)
@@ 6 Numerical Results
To evaluate the effectiveness of RCPI a series of examples were solved using a range of algorithms including PAI, Hoffman Karp (HK), Filar Tolwinski’s algorithm (FT), robust value iteration (VI), a variation of Hoffman Karp (PPI) (Ho2021), a variation on PAI (WS) (Winnicki2023), and our algorithm (RCPI). To simplify comparison, RCPI’s hyperparameter $m$ was either set to 0, producing a method that never fixed its evaluation step, called RCPI0, or $m$ was left unbounded, making a method that always fixed its evaluation step, called RCPI. The full source code for the algorithms and domains is available at https://github.com/keithbadger/Fast-Policy-Iteration-for-Markov-Games-and-Robust-MDPs.
For each domain, there was a smaller set of problems solved with $\gamma$ set to $0.5$, $0.75$, $0.9$, and $0.99$, and a larger set of problems where we set $\gamma$ to $0.9$, $0.99$, and $0.999$. We use the smaller problems in order to evaluate the slower algorithms in a reasonable time. The larger problems can only be solved by the faster methods within our time limits. We allocated a larger time budget to VI to establish a reference.
@@
Several domains were tested for robust MDPs including gamblers ruin (Kallenberg2022), gridworld (Sutton1998, section 6.5), and inventory management (Puterman2005, section 3.2). From Table˜1, each algorithm maintained the same relative performance from Markov games, except that WS and VI did comparatively better in the gambler’s ruin. WS and VI did well because the optimal betting scheme in the non-robust version of gambler’s ruin is to bet \$1 if the win rate is greater than $50\%$ and to bet all money otherwise. The gambler repeats this action until reaching the maximum capital, when it obtains the reward. The result is a singular optimal state trajectory following the current state. The reward from obtaining the maximum capital does not affect the current state’s policy until there is a Bellman backup for every state in the optimal trajectory following it. This type of domain favors methods with inexpensive evaluations, as only states that reward has been reached via policy improvement will be worth evaluating. Time spent doing exact evaluations for the other states does not provide any benefit.
@@ 7 Conclusion
It remains to be seen if there is a way to fix the algorithm of Filar and Tolwinski that keeps the Newtonian interpretation of PAI with $\psi_{2}(\bm{v})^{2}$ as the objective function. Such a solution could provide insight into how to deal with discontinuities caused by the min and max operations more generally. It is not known how to optimize the hyperparameter of RCPI $m$ which could reduce the level of knowledge required to use RCPI effectively.
Acknowledgments
We thank the anonymous reviewers for their detailed reviews and thoughtful comments, which significantly improved the paper’s clarity. This work was supported, in part, by NSF grants 2144601 and 2218063 and ONR grant N0001425GI01179.
@@ A.2 Definitions for Robust MDPs
In this section, we summarize the pertinent definitions of value functions and Bellman operators in Robust MDPs. The value function $\bm{v}^{\bm{\pi},\bm{p}}\in\mathbb{R}^{S}$ is defined as for each $\bm{\pi}\in\Pi$ and $\bm{p}\in\mathcal{P}$ as (Wiesemann2013; Ho2022):
$v^{\bm{\pi},\bm{p}}_{s}\;:=\;\mathbb{E}_{\bm{\pi},\bm{p}}^{s}\left[\sum_{t=0}^{\infty}\gamma^{t}r(\tilde{s}_{t},\tilde{a}_{t})\right],\quad\forall s\in\mathcal{S}.$ (14)
@@
$P^{\bm{\pi},\bm{p}}_{s,s^{\prime}}:=\sum_{a\in\mathcal{A}}\pi_{a}(s)\cdot p_{s^{\prime}}(s,a),\qquad r^{\bm{\pi}}_{s}:=\sum_{a\in\mathcal{A}}\pi_{a}(s)\cdot r(s,a).$
@@
The Bellman operators $\mathfrak{T}^{\bm{\pi},\bm{p}}$ and $\mathfrak{T}^{\star}$ are monotone and $\gamma$-contractive in the $L_{\infty}$ norm (Kallenberg2022, theorem 10.5). They can be computed by solving a linear program (Ho2022, appendix C) or by a specialized algorithm (Ho2022, section 6).
@@ 4   Select $(\bm{\pi}^{k},\bm{\sigma}^{k})\in\mathfrak{B}^{\star}\bm{v}^{k-1}$;
Consider the policy such that $\bm{\sigma}^{1}(s_{1})=[1,0]$. Then, Algorithm˜1 computes the step direction as:
@@
To satisfy the starting constraint given from (Filar1991, theorem 3.3) that $\bm{v}^{0}=r_{\max}\cdot\bm{1}$ where $r_{\max}=\max_{a\in\mathcal{A},b\in\mathcal{B},s^{\prime}\in\mathcal{S}}|r(a,b,s^{\prime})|$, consider a MG with the transition probabilities and rewards defined in Figure˜4 and the discount factor $\gamma=0.8$.
@@
which is always positive for $\alpha\in(0,1]$, thus using the initial value from (Filar1991, theorem 3.3) does not fix FT’s convergence issues.
@@ Appendix C Additional Material for Section 5
This linear program contains $A+B$ constraints and $A$ variables, requiring $O\left((A+B)^{1.5}A^{2}\log(\delta^{-1})\right)$ arithmetic operations to solve to a precision of $\delta$ (InteriorPointMethods, section 10.1). Therefore, the total cost of computing $\mathfrak{B}^{\delta}\bm{v}$ and $\mathfrak{T}^{\delta}\bm{v}$ is the same as generating $\bm{G}^{\bm{v},s}$ and solving the linear program (C.1) for every state, producing the bound in (9). ∎